OpenStudy (lxelle):

Melons are sold in three sizes: small, medium and large. The weights follow a normal distribution with mean 450 grams and standard deviation 120 grams. Melons weighing less than 350 grams are classified as small. (i) Find the proportion of melons which are classified as small. (ii) The rest of the melons are divided in equal proportions between medium and large. Find the weight above which melons are classified as large.

OpenStudy (lxelle):

I've gotten the first answer as 20.2%

OpenStudy (anonymous):

what r the choices u can choose from

OpenStudy (dangerousjesse):

These types of questions don't normally have choices :)

OpenStudy (lxelle):

@dillon754 wtf you're crapping ?

OpenStudy (anonymous):

what

OpenStudy (dangerousjesse):

Once you hit Algebra I, you stop having multiple choice :)

OpenStudy (lxelle):

This is statistics.

OpenStudy (anonymous):

that sucks

OpenStudy (lxelle):

Get lost if you're not helping here. Thanks.

OpenStudy (anonymous):

im not good at algebra im really sorry

OpenStudy (dangerousjesse):

For the first one, it's \(\large 20.33%\). Please be nice :)

OpenStudy (dangerousjesse):

For the second one, it's 480 grams.

OpenStudy (amistre64):

and how do we get to 480 as a solution?

OpenStudy (dangerousjesse):

\(\huge1 - 0.2033 = 0.7967 \) \(\huge \frac{0.7967}{2} = 0.39835 \)

OpenStudy (dangerousjesse):

So the proportion of large melons would be 0.39835.

OpenStudy (amistre64):

seems like a fair approach :)

OpenStudy (anonymous):

im really sorry

OpenStudy (dangerousjesse):

The x value which separates the top 0.39835 melons from the rest is about \(+0.25\). The weight of a large melon = y = the mean \(+ (z \times standard~deviation)\)

OpenStudy (lxelle):

I don't understand . :(

OpenStudy (dangerousjesse):

So you have \(450+(0.25\times 120)=?\)

OpenStudy (dangerousjesse):

What don't you understand? :)

OpenStudy (lxelle):

Slow down. I don't get your second last comment.

OpenStudy (lxelle):

What do you mean by the x value separates and all.

OpenStudy (dangerousjesse):

I meant x, not z, sorry. The proportion of the melons that are large compared to those that are medium is 39.835%. When we average in the difference between large and small, we get about 25%.

OpenStudy (dangerousjesse):

Since the average weight of a melon is 450 grams, we multiply 0.25 and the standard deviation (120) together and add the product to 450, which gives us 480 grams as the size of the smallest large melon :)

OpenStudy (lxelle):

I still don't get it, :(

OpenStudy (lxelle):

How did you get 25%?

OpenStudy (lxelle):

You just freaking copied whatever the yahoo answer said.

OpenStudy (dangerousjesse):

25% is the average differentiation between large and small, and large and medium. Pardon me?

OpenStudy (amistre64):

you found the percentage of small, the information tells us that med and large split the difference.

OpenStudy (amistre64):

small + (1-small)/2 is the dividing line between med and large

OpenStudy (amistre64):

assuming you want a left tail that is

OpenStudy (amistre64):

when we know the percentage that is the dividing line, we can use an inverse normal function to determine a z score, and thereby find the weight that is associated with it

OpenStudy (lxelle):

How to get. 0.25??

OpenStudy (lxelle):

Is it possible to illustrate the graph for me?

OpenStudy (amistre64):

i cant draw a the moment, just have a keyboard the percentages are divided up along the distribution like this: k% n% n% |-------------|---------------|---------------| small med large such that k% + n% + n% = 100%, agreed?

OpenStudy (lxelle):

Yes!

OpenStudy (amistre64):

we know the zscore for small is: (350-450)/120 = -0.8333, which we can table or calculator as; 20.23 % agreed?

OpenStudy (lxelle):

Yup.

OpenStudy (amistre64):

so: 20.23% + 2n% = 100% n% = (79.77%)/2 = 39.885% this tells us the split between med and large is 20.23% + 39.885% = 60.115% agreed?

OpenStudy (lxelle):

Yes

OpenStudy (amistre64):

now, we need to work an inverse normal function, it takes a percentage and gives us a zscore. we can either do this by a table finding the closest field value to .60115 or a calculator that has an invnorm function. the wolf gives us: invnorm(.60115) = 0.256 for the associated zscore.

OpenStudy (amistre64):

we can now use this inn the zscore formula to find x:\[z=\frac{x-mean}{sd}\] solving for x we get: \[mean + z(sd)=x\] using our values we have:\[450+0.256(120) = x\]which is the dividing line between med and large

OpenStudy (lxelle):

Thanks so much. I finally understand. :)